a cyclist and their bike have a combined mass of 94kg and a kinetic energy of 2303J. what is the speed of the cyclist?
1
Expert's answer
2013-06-28T07:20:38-0400
As we know, kinetic energy is defined by the following formula:
Ek=mv^2/2
where m-mass, v-velocity


whence magnitude of velocity (which is called speed) equals v=sqrt(2*Ek/m)
So in our case we have:
v=sqrt(2*2303/94)=7 m/s